OpenStudy (anonymous):

If f(x) = 5x-4 and g(x) =6x+3, what is f(g(x))? 1) 30x-21 2) 30x-1 3) 30x+11 4) 30x+12

OpenStudy (anonymous):

30x+11

OpenStudy (anonymous):

3

OpenStudy (anonymous):

f(g(x))=f(6x+3)=5(6x+3)-4=30x+11